A refinery produces two types of lubricating oil, type A and type B. To produce 1000 litres of type A oil requires two hoursType or paste question here
Answer #1

(a) The Decision Variables are:

  • Type A Production per day
  • Type B Production per day

(b) The Objective function is to Maximize the Profit, where

  • Profit = Type A Profit + Type B Profit
  • Type A Profit = 40 * (Type A Production / 1000)   
  • Type B Profit = 50 * (Type B Production / 1000)  

(c) subject to the following Constraints:

  • Usage of P1 plant = ((2 * Type A Prodn / 1000) + (4 * Type B Prodn / 1000)) <= 12 hrs per day
  • Usage of P2 plant = ((2 * Type A Prodn / 1000) + (2 * Type B Prodn / 1000)) <= 7 hrs per day
  • Usage of P3 plant = ((6 * Type A Prodn / 1000) + (3 * Type B Prodn / 1000)) <= 20    hrs per day
  • Type A Production >= 0
  • Type B Production >= 0

(d) Feasible Region with below corner points and objective function:

Corner Points OF
0, 0 0
0, 3000 150
3333, 0 133.32
1000, 2500 165
3166.67, 333.33 143.32